# Constrained free energy minimization for the design of thermal states and stabilizer thermodynamic systems

### v2

@@ Abstract
A quantum thermodynamic system is described by a Hamiltonian and a list of conserved, non-commuting charges, and a fundamental goal is to determine the minimum energy of the system subject to constraints on the charges. Recently, [Liu et al., arXiv:2505.04514] proposed first- and second-order classical and hybrid quantum-classical algorithms for solving a dual chemical potential maximization problem, and they proved that these algorithms converge to global optima by means of gradient-ascent approaches. In this paper, we benchmark these algorithms on several problems of interest in thermodynamics, including one- and two-dimensional quantum Heisenberg models with nearest and next-to-nearest neighbor interactions and with the charges set to the total $x$, $y$, and $z$ magnetizations. We also offer an alternative compelling interpretation of these algorithms as methods for designing ground and thermal states of controllable Hamiltonians, with potential applications in molecular and material design. Furthermore, we introduce stabilizer thermodynamic systems as thermodynamic systems based on stabilizer codes, with the Hamiltonian constructed from a given code’s stabilizer operators and the charges constructed from the code’s logical operators. We benchmark the aforementioned algorithms on several examples of stabilizer thermodynamic systems, including those constructed from the one-to-three-qubit repetition code, the perfect one-to-five-qubit code, and the two-to-four-qubit error-detecting code. Finally, we observe that the aforementioned hybrid quantum-classical algorithms, when applied to stabilizer thermodynamic systems, can serve as alternative methods for encoding quantum information into stabilizer codes at a fixed temperature, and we provide an effective method for warm-starting these encoding algorithms whenever a single qubit is encoded into multiple physical qubits.
@@ I.1 Background and motivation
Quantum computation has the potential to cause a paradigm shift in several areas of computer science. After the initial proposals of quantum speedups for factoring [1] and unstructured search [2], it was postulated that quantum computation could offer speedups for optimization, with applications to physics, combinatorial optimization, machine learning, and beyond (see, e.g., [3] for a recent review). While this possibility still remains the subject of ongoing research, concrete quantum algorithms for solving semi-definite optimization problems have been proposed starting nearly a decade ago [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Some of these algorithms have guaranteed runtimes [4, 5, 6, 7, 8, 10, 11, 16], while others are heuristic in nature [9, 12, 13, 14, 15], falling under the umbrella of variational quantum algorithms [17].
@@
Liu et al. recently identified a concrete link between quantum thermodynamics and semi-definite optimization [16], having the benefit of bridging two previously unlinked foundational areas and bringing insights from quantum thermodynamics into the analysis of semi-definite optimization. As identified in [16], a special class of semi-definite optimization problems, known as constrained energy minimization problems, are relevant in quantum thermodynamics. In particular, these problems are defined in terms of a Hamiltonian and a list of conserved, non-commuting charges, with the latter imposing constraints on the state of the physical system. The goal then is to minimize the energy of the system, while respecting the imposed constraints. Generally speaking, determining the minimum energy of a quantum physical system is a primary goal in physics [18], typically being the first step employed in computing energetic properties of molecules and materials, and thus having applications in materials science [19], condensed-matter physics [20], and quantum chemistry [21]. Going beyond this application, and as discussed in this paper, constrained energy minimization has applications in the design of molecules and materials having desired properties.
@@
Beyond establishing a link between quantum thermodynamics and semi-definite optimization, Ref. [16] proposed various classical and hybrid quantum-classical (HQC) algorithms for solving semi-definite programs, being motivated by physical intuition coming from quantum thermodynamics when applied to the special class of constrained energy minimization problems. These algorithms, hereafter referred to as the LMPW algorithms, have provable guarantees for their runtimes and are based on the observation that parameterized thermal states, also known as non-Abelian thermal states [22, 23, 24] or quantum Boltzmann machines [25, 26, 27], are optimal for solving constrained free-energy minimization problems, which closely approximate the original constrained energy minimization problems for sufficiently low temperatures. Indeed, Ref. [16] established that the optimization landscapes of dual chemical potential maximization problems are concave in the parameters of the thermal states, so that conventional optimization techniques like gradient ascent, stochastic gradient ascent, and their variants [28] are guaranteed to converge. The HQC algorithms are variational in nature, using a quantum computer for two purposes only, which include the preparation of parameterized thermal states and estimating expectations of observables (i.e., the Hamiltonian and non-commuting charges). A classical computer is then employed for storing and updating the parameters of the parameterized thermal states as the HQC algorithm progresses toward convergence. The LMPW HQC algorithms make use of first-derivative information to determine the next step in a search and can additionally incorporate second-derivative information to improve their convergence to globally optimal solutions.
@@ I.2 Summary of contributions
In this paper, we apply the LMPW algorithms to concrete examples of constrained energy minimization problems of interest in quantum thermodynamics, including the one- and two-dimensional Heisenberg models [37] with nearest and next-to-nearest neighbor interactions, and with the non-commuting charges set to the total magnetizations in the $x$, $y$, and $z$ directions. Such models describe interacting spin systems on a lattice and are essential for understanding magnetic materials [38].
@@
As a key theoretical contribution, our paper also establishes a bridge between quantum error correction [39, 40, 41, 42, 43] and conserved, non-commuting charges in quantum thermodynamics, by observing that every stabilizer code [44, 45] can be understood as corresponding to a thermodynamic system with conserved, non-commuting charges. In particular, we set the Hamiltonian to be in correspondence with the stabilizer operators of the code, while the non-commuting charges are in correspondence with the logical operators of the code. We refer to such a thermodynamic system as a stabilizer thermodynamic system. Let us note that stabilizer Hamiltonians were proposed in [46] and have been investigated substantially since then (see, e.g., [47, 48, 49, 50, 51, 52, 53, 54, 55]). However, interpreting stabilizer and logical operators in terms of Hamiltonians and conserved, non-commuting charges, respectively, in a thermodynamic system has not been considered previously, to the best of our knowledge.
Moreover, we utilize this correspondence to demonstrate that the LMPW HQC algorithms can serve as thermodynamics-inspired HQC algorithms for encoding quantum information into quantum error-correcting codes, which is rather different from the standard approach that employs an encoding circuit to do so (see, e.g., [56, Chapter 4] and [57, Section 3.5]). The approach is instead based on constrained energy minimization and the fact that the LMPW HQC algorithms, as applied in this context, prepare low-temperature thermal states of stabilizer thermodynamic systems. We suspect that this approach will be useful in experimental scenarios for which it is accessible to prepare low-temperature thermal states of stabilizer thermodynamic systems.
@@ I.3 Overview of numerical simulations
Before proceeding with the details of our paper, let us first elaborate on some aspects of the numerical simulations that we conducted. Given that all of the constrained energy minimization problems we consider can be formulated as semi-definite programs (SDPs) [63, Section 4.6.2], the performance of the LMPW algorithms can be benchmarked against the values returned by standard SDP solvers. These solvers are based on the interior-point method [64, 65, 66] and are so reliable that we can view their outputs as essentially the true optimal values of the constrained energy minimization problems. We test out four variants of the LMPW algorithms, which include a first-order classical algorithm, a second-order classical algorithm, a first-order HQC algorithm, and a second-order HQC algorithm. Considering the classical algorithms in addition to the HQC algorithms is useful, as the classical algorithms can be understood as fully noiseless simulators of the HQC algorithms, such that the only kind of error involved when using them is the algorithmic error and not any other kind of error, e.g., shot noise arising from estimating expectations of observables. For this reason, we can benchmark the HQC algorithms against the classical algorithms, in addition to the true values generated by standard SDP solvers.
@@ II.2.1 Review of constrained energy minimization
This form of the gradient allows for calculation by means of a classical algorithm for small system sizes or for estimation by means of quantum algorithm equipped with subroutines for thermal state preparation of $\rho_{T}(\mu)$ and measurement of each observable $Q_{i}$.
@@ II.2.2 LMPW first-order classical algorithm
We have phrased the LMPW classical algorithm slightly differently from how it was presented previously in [16, Algorithm 1], in order to simplify its presentation. The main difference is the introduction of the stopping parameter $\delta$ above, as a simple way of ending the algorithm after convergence has been achieved (i.e., when the norm of the gradient is small, further changes to $\mu$ are small and convergence near a globally optimal solution is guaranteed at this point by [77, Corollary 3.5]).
@@ II.2.3 LMPW first-order hybrid quantum-classical algorithm
As presented above, Algorithm 2 is a simplification of the first-order HQC algorithm proposed in [16, Algorithm 2]. Since the gradient calculations are replaced by gradient estimates, the latter are random variables, and the algorithm thus employs stochastic gradient ascent in its search for a globally optimal solution (see, e.g., [77, Chapter 5] and [28, Section 6]). Since one of the goals of [16] was to prove that the algorithm converges to a global optimum, the presentation of the algorithm was more involved in order to provide a theoretical guarantee on its performance. However, in practice, we find in our simulations that the above simplification of the algorithm works well for converging near the global optimum.
@@ II.2.4 LMPW second-order classical algorithm
As noted in [16, Lemmas 6 and 8], the Hessian is negative semi-definite, implying that the objective function $f(\mu)$ in (16) is concave. Let us denote the Hessian matrix evaluated at $\nu\in\mathbb{R}^{c}$ by $\nabla_{\mu}^{2}f(\nu)$, so that it has matrix elements
@@
Due to the Hessian being negative semi-definite, the minus sign in (27) effectively cancels out with the implicit minus sign from the Hessian $\nabla_{\mu}^{2}f(\mu^{m-1})$, so that each gradient update always proceeds in a positive direction and corresponds to an ascent procedure.
@@ II.2.5 LMPW second-order hybrid quantum-classical algorithm
Finally, let us recall a second-order HQC algorithm for constrained energy minimization, as proposed in [16, Section VI]. A key observation, originally made in the context of a general-purpose natural gradient algorithm for quantum Boltzmann machines [82, Theorem 2], is that the Hessian matrix in (24) admits the following alternative expression [16, Lemma 6]:
@@
and $p(t)$ is the following high-peak tent probability density defined for $t\in\mathbb{R}$ [86, Eq. (9)]:
@@
In addition to the backtracking approach mentioned after Algorithm 3, we also regularize the estimate of the Hessian matrix, by adding a scaled identity matrix to it, in order to ensure that it is negative semi-definite. By employing both backtracking and regularization, we find in our simulations that the algorithm converges in all cases. Without doing so, the estimate of the Hessian matrix need not be negative semi-definite, due to noise inherent in the sampling process, and the algorithm need not converge in such a case.
@@ II.3 Quantum Heisenberg models
Special cases of the general model presented in (40)–(46) include the one- and two-dimensional quantum Heisenberg models with nearest-neighbor and next-to-nearest-neighbor interactions. In the case of a one-dimensional model, the qubits are arranged on a line with edges placed between nearest neighbors, for the nearest neighbor case, and edges additionally placed between next-to-nearest neighbors, for the next-to-nearest neighbor case. To simplify the model further, we set the coupling coefficients to be a fixed value $J$ between nearest neighbors, and for next-to-nearest neighbors, we set the coupling coefficients to be fixed at a constant value, say, $\lambda J$, where $\lambda\in[0,1]$. For a two-dimensional model, the main difference is that the qubits are arranged on a square lattice, and we take similar conventions, as mentioned above, for the nearest-neighbor and next-to-nearest neighbor cases. See Figure 1 for a depiction of these models.
Figure 1: Depiction of one- and two-dimensional Heisenberg models. (Top) Qubits of the one-dimensional model are arranged on a line with edges placed between nearest neighbors (black lines), and edges additionally placed between next-to-nearest neighbors (blue curves). (Bottom) Qubits of the two-dimensional model are arranged on a square lattice, with nearest neighbor and next-to-nearest neighbor interactions shown in black and blue lines, respectively.
@@ II.4 Stabilizer codes for quantum error correction
where $L\mathcal{S}$ is the group realized by left multiplying all elements of $\mathcal{S}$ by $L$. For a stabilizer code, the normalizer $N(\mathcal{S})$ of $\mathcal{S}$ coincides with its centralizer [56, Section 3.2]:

### v3

@@ Abstract
A quantum thermodynamic system is described by a Hamiltonian and a list of conserved, non-commuting charges, and a fundamental goal is to determine the minimum energy of the system subject to constraints on the charges. Recently, [Liu et al., arXiv:2505.04514] proposed first- and second-order classical and hybrid quantum-classical algorithms for solving a dual chemical potential maximization problem, and they proved that these algorithms converge to global optima by means of gradient-ascent approaches. In this paper, we benchmark these algorithms on several problems of interest in thermodynamics, including one- and two-dimensional quantum Heisenberg models with nearest- and next-nearest neighbor interactions and with the charges set to the total $x$, $y$, and $z$ magnetizations. We also offer an alternative compelling interpretation of these algorithms as methods for designing ground and thermal states of controllable Hamiltonians, with potential applications in molecular and material design. Furthermore, we introduce stabilizer thermodynamic systems as thermodynamic systems based on stabilizer codes, with the Hamiltonian constructed from a given code’s stabilizer operators and the charges constructed from the code’s logical operators. We benchmark the aforementioned algorithms on several examples of stabilizer thermodynamic systems, including those constructed from the one-to-three-qubit repetition code, the perfect one-to-five-qubit code, and the two-to-four-qubit error-detecting code. Finally, we observe that the aforementioned hybrid quantum-classical algorithms, when applied to stabilizer thermodynamic systems, can serve as alternative methods for encoding quantum information into stabilizer codes at a fixed temperature, and we provide an effective method for warm-starting these encoding algorithms whenever a single qubit is encoded into multiple physical qubits.
@@ I.1 Background and motivation
Quantum computation has the potential to cause a paradigm shift in several areas of computer science. After the initial proposals of quantum speedups for factoring [1] and unstructured search [2], it was postulated that quantum computation could offer speedups for optimization, with applications to physics, combinatorial optimization, machine learning, and beyond (see, e.g., [3] for a recent review). While this possibility still remains the subject of ongoing research, concrete quantum algorithms for solving semidefinite optimization problems have been proposed starting nearly a decade ago [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Some of these algorithms have guaranteed runtimes [4, 5, 6, 7, 8, 10, 11, 16], while others are heuristic in nature [9, 12, 13, 14, 15], falling under the umbrella of variational quantum algorithms [17].
@@
Liu et al. recently identified a concrete link between quantum thermodynamics and semidefinite optimization [16], having the benefit of bridging two previously unlinked foundational areas and bringing insights from quantum thermodynamics into the analysis of semidefinite optimization. As identified in [16], a special class of semidefinite optimization problems, known as constrained energy minimization problems, are relevant in quantum thermodynamics. In particular, these problems are defined in terms of a Hamiltonian and a list of conserved, non-commuting charges, with the latter imposing constraints on the state of the physical system. The goal then is to minimize the energy of the system, while respecting the imposed constraints.
Generally speaking, determining the minimum energy of a quantum physical system is a primary goal in physics [18], typically being the first step employed in computing energetic properties of molecules and materials, and thus having applications in materials science [19], condensed-matter physics [20], and quantum chemistry [21]. Going beyond this application, and as discussed in this paper, constrained energy minimization has applications in the design of molecules and materials having desired properties. In particular, the ability to enforce constraints on conserved quantities (such as particle number or spin) allows for the precise targeting of specific electronic symmetry sectors – a capability that is often challenging for standard unconstrained variational approaches which may collapse to incorrect sectors (e.g., neutral species instead of ions) [22, 23, 24].
@@
Beyond establishing a link between quantum thermodynamics and semidefinite optimization, Ref. [16] proposed various classical and hybrid quantum-classical (HQC) algorithms for solving semidefinite programs, being motivated by physical intuition coming from quantum thermodynamics when applied to the special class of constrained energy minimization problems. These algorithms, hereafter referred to as the LMPW algorithms, have provable guarantees for their runtimes and are based on the observation that parameterized thermal states, also known as non-Abelian thermal states [25, 26, 27] or quantum Boltzmann machines [28, 29, 30], are optimal for solving constrained free-energy minimization problems, which closely approximate the original constrained energy minimization problems for sufficiently low temperatures. Indeed, Ref. [16] established that the optimization landscapes of dual chemical potential maximization problems are concave in the parameters of the thermal states, so that conventional optimization techniques like gradient ascent, stochastic gradient ascent, and their variants [31] are guaranteed to converge. The HQC algorithms are variational in nature, using a quantum computer for two purposes only, which include the preparation of parameterized thermal states and estimating expectations of observables (i.e., the Hamiltonian and non-commuting charges). A classical computer is then employed for storing and updating the parameters of the parameterized thermal states as the HQC algorithm progresses toward convergence. The LMPW HQC algorithms make use of first-derivative information to determine the next step in a search and can additionally incorporate second-derivative information to improve their convergence to globally optimal solutions.
@@
This capability is particularly timely, as recent developments in quantum thermodynamics have highlighted the rich physics of systems with conserved, non-commuting charges [40, 27]. Unlike standard commuting charges, non-commutation relations can fundamentally alter many-body phenomena, leading to effects such as the modification of entropy production [41] and the enhancement of entanglement entropy [42]. Furthermore, these charges play a critical role in thermalization dynamics; for instance, they can invalidate the standard Eigenstate Thermalization Hypothesis (ETH) [43] or prevent certain forms of thermalization entirely [44]. However, while the dynamical consequences of non-commuting charges have been actively explored, the rigorous construction and calculation of the resulting equilibrium states remains a computational challenge. Our work addresses this gap by providing a suite of algorithms to explicitly construct these states via free energy minimization, offering a tool to verify the equilibrium endpoints predicted by these thermodynamic theories.
@@ I.2 Summary of contributions
In this paper, we apply the LMPW algorithms to concrete examples of constrained energy minimization problems of interest in quantum thermodynamics, including the one- and two-dimensional Heisenberg models [45] with nearest and next-nearest neighbor interactions, and with the non-commuting charges set to the total magnetizations in the $x$, $y$, and $z$ directions. Such models describe interacting spin systems on a lattice and are essential for understanding magnetic materials [46].
@@
As a key theoretical contribution, our paper also establishes a bridge between quantum error correction [47, 48, 49, 50, 51] and conserved, non-commuting charges in quantum thermodynamics, by observing that every stabilizer code [52, 53] can be understood as corresponding to a thermodynamic system with conserved, non-commuting charges. In particular, we set the Hamiltonian to be in correspondence with the stabilizer operators of the code, while the non-commuting charges are in correspondence with the logical operators of the code. We refer to such a thermodynamic system as a stabilizer thermodynamic system.
Let us note that stabilizer Hamiltonians were proposed in [54] and have been investigated substantially since then (see, e.g., [55, 56, 57, 58, 59, 60, 61, 62, 63]). However, interpreting logical operators in terms of conserved, non-commuting charges in a thermodynamic system has not been considered previously, to the best of our knowledge. While the aforementioned previous works largely studied the passive properties of stabilizer Hamiltonians (e.g., thermal stability and memory lifetime [56, 64, 65, 63]), our work utilizes stabilizer thermodynamic systems for active state preparation. By formally treating the logical operators as non-commuting charges with associated chemical potentials within a thermodynamic ensemble, we demonstrate that the LMPW HQC algorithms can serve as thermodynamics-inspired HQC algorithms for encoding quantum information into quantum error-correcting codes, which is rather different from the standard approach that employs an encoding circuit to do so (see, e.g., [66, Chapter 4], [67], and [68, Section 3.5]). The approach is instead based on constrained energy minimization and the fact that the LMPW HQC algorithms, as applied in this context, prepare low-temperature thermal states of stabilizer thermodynamic systems. We envision that this approach will be useful in experimental scenarios for which it is accessible to prepare low-temperature thermal states of stabilizer thermodynamic systems.
@@ I.3 Overview of numerical simulations
Before proceeding with the details of our paper, let us first elaborate on some aspects of the numerical simulations that we conducted. Given that all of the constrained energy minimization problems we consider can be formulated as semidefinite programs (SDPs) [74, Section 4.6.2], the performance of the LMPW algorithms can be benchmarked against the values returned by standard SDP solvers. These solvers are based on the interior-point method [75, 76, 77] and are so reliable that we can view their outputs as essentially the true optimal values of the constrained energy minimization problems. We test out four variants of the LMPW algorithms, which include a first-order classical algorithm, a second-order classical algorithm, a first-order HQC algorithm, and a second-order HQC algorithm. Considering the classical algorithms in addition to the HQC algorithms is useful, as the classical algorithms can be understood as fully noiseless simulators of the HQC algorithms, such that the only kind of error involved when using them is the algorithmic error and not any other kind of error, e.g., shot noise arising from estimating expectations of observables. For this reason, we can benchmark the HQC algorithms against the classical algorithms, in addition to the true values generated by standard SDP solvers.
@@ II.2.1 Review of constrained energy minimization
This form of the gradient allows for calculation by means of a classical algorithm for small system sizes or for estimation by means of a quantum algorithm equipped with subroutines for thermal state preparation of $\rho_{T}(\mu)$ and measurement of each observable $Q_{i}$.
@@ II.2.2 LMPW first-order classical algorithm
We have phrased the LMPW classical algorithm slightly differently from how it was presented previously in [16, Algorithm 1], in order to simplify its presentation. The main difference is the introduction of the stopping parameter $\delta$ above, as a simple way of ending the algorithm after convergence has been achieved (i.e., when the norm of the gradient is small, further changes to $\mu$ are small and convergence near a globally optimal solution is guaranteed at this point by [88, Corollary 3.5]).
@@ II.2.3 LMPW first-order hybrid quantum-classical algorithm
As presented above, Algorithm 2 is a simplification of the first-order HQC algorithm proposed in [16, Algorithm 2]. Since the gradient calculations are replaced by gradient estimates, the latter are random variables, and the algorithm thus employs stochastic gradient ascent in its search for a globally optimal solution (see, e.g., [88, Chapter 5] and [31, Section 6]). Since one of the goals of [16] was to prove that the algorithm converges to a global optimum, the presentation of the algorithm was more involved in order to provide a theoretical guarantee on its performance. However, in practice, we find in our simulations that the above simplification of the algorithm works well for converging near the global optimum.
@@ II.2.4 LMPW second-order classical algorithm
As noted in [16, Lemmas 6 and 8], the Hessian is negative semidefinite, implying that the objective function $f(\mu)$ in (16) is concave. Let us denote the Hessian matrix evaluated at $\nu\in\mathbb{R}^{c}$ by $\nabla_{\mu}^{2}f(\nu)$, so that it has matrix elements
@@
Due to the Hessian being negative semidefinite, the minus sign in (27) effectively cancels out with the implicit minus sign from the Hessian $\nabla_{\mu}^{2}f(\mu^{m-1})$, so that each gradient update always proceeds in a positive direction and corresponds to an ascent procedure.
@@ II.2.5 LMPW second-order hybrid quantum-classical algorithm
Finally, let us recall a second-order HQC algorithm for constrained energy minimization, as proposed in [16, Section VI]. A key observation, originally made in the context of a general-purpose natural gradient algorithm for quantum Boltzmann machines [93, Theorem 2], is that the Hessian matrix in (24) admits the following alternative expression [16, Lemma 6]:
@@
and $p(t)$ is the following high-peak tent probability density defined for $t\in\mathbb{R}$ [97, Eq. (9)]:
@@
In addition to the backtracking approach mentioned after Algorithm 3, we also regularize the estimate of the Hessian matrix, by adding a scaled identity matrix to it, in order to ensure that it is negative semidefinite. By employing both backtracking and regularization, we find in our simulations that the algorithm converges in all cases. Without doing so, the estimate of the Hessian matrix need not be negative semidefinite, due to noise inherent in the sampling process, and the algorithm need not converge in such a case.
@@ II.3 Quantum Heisenberg models
Special cases of the general model presented in (40)–(46) include the one- and two-dimensional quantum Heisenberg models with nearest-neighbor and next-nearest-neighbor interactions. In the case of a one-dimensional model, the qubits are arranged on a line with edges placed between nearest neighbors, for the nearest neighbor case, and edges additionally placed between next-nearest neighbors, for the next-nearest neighbor case. To simplify the model further, we set the coupling coefficients to be a fixed value $J$ between nearest neighbors, and for next-nearest neighbors, we set the coupling coefficients to be fixed at a constant value, say, $\lambda J$, where $\lambda\in[0,1]$. For a two-dimensional model, the main difference is that the qubits are arranged on a square lattice, and we take similar conventions, as mentioned above, for the nearest-neighbor and next-nearest-neighbor cases. See Figure 1 for a depiction of these models.
Figure 1: Depiction of one- and two-dimensional Heisenberg models. (Top) Qubits of the one-dimensional model are arranged on a line, with edges placed between nearest neighbors (black lines) and additionally between next-nearest neighbors (blue curves). (Bottom) Qubits of the two-dimensional model are arranged on a square lattice, with nearest-neighbor and next-nearest-neighbor interactions shown in black and blue lines, respectively.
@@ II.4 Stabilizer codes for quantum error correction
where $L\mathcal{S}$ is the group realized by left multiplying all elements of $\mathcal{S}$ by $L$. For a stabilizer code, the normalizer $N(\mathcal{S})$ of $\mathcal{S}$ coincides with its centralizer [66, Section 3.2]: